OpenStudy (anonymous):

find the 1st 5 perfect squares of x such that 2 times the square root of x is also a perfect square

OpenStudy (mathmate):

If I interpret correctly, then the you look for x^2 such that 2sqrt(x)=perfect square. This implies: 2sqrt(x)=y^2 where x, y are integers. square both sides and divide by 4: x=y^4/4 Put y=0, x=0 y=1, x is not an integer y=2, x=4 ... Can you complete the solution?

OpenStudy (anonymous):

thanks

OpenStudy (mathmate):

You're welcome! :)